Pronghorn Appliance Center is an experienced home appliance dealer. Pronghorn Appliance Center also offers a number of services for the home appliances that it sells. Assume that Pronghorn Appliance Center sells ovens on a standalone basis. Pronghorn Appliance Center also sells installation services and maintenance services for ovens. However, Pronghorn Appliance Center does not offer installation or maintenance services to customers who buy ovens from other vendors. Pricing for ovens is as follows. Oven only $800 Oven with installation service 850 Oven with maintenance services 990 Oven with installation and maintenance services 1,000 In each instance in which maintenance services are provided, the maintenance service is separately priced within the arrangement at $190. Additionally, the incremental amount charged by Pronghorn Appliance Center for installation approximates the amount charged by independent third parties. Ovens are sold subject to a general right of return. If a customer purchases an oven with installation and/or maintenance services, in the event Pronghorn Appliance Center does not complete the service satisfactorily, the customer is only entitled to a refund of the portion of the fee that exceeds $800.

Assume that a customer purchases an oven with both installation and maintenance services for $1,000.

Indicate the amount of revenue that should be allocated to the oven, the installation, and to the maintenance contract
Answer #1
Revenue allocated:
Fair value % of total fair value Total sales value Revenue allocated
a b a*b
Oven 800 76.92% 1000 769
Installation (Note:1) 50 4.81% 1000 48
Maintenance (Note:1) 190 18.27% 1000 183
1040 1000
Note:1
Stand alone price of installation=Oven with installation price-Price of oven=850-800=$ 50
Stand alone price of maintenance=Oven with mainteance service-Price of oven=990-800=$ 190